OpenStudy (anonymous):

xy' - y = x^3 in the method of variation of parameters how can i solve?

OpenStudy (anonymous):

oh this one is not as fun haha. Anyways do you know Kramers rule?

OpenStudy (anonymous):

yeah i know

OpenStudy (anonymous):

alright so first find your complementary using an auxilary equation ( putting the equation in terms of m)

OpenStudy (anonymous):

i learnt not to much but in the examples it does not have x in the LHS

OpenStudy (anonymous):

actually in this case you will have to use another way other than auxilary

OpenStudy (anonymous):

divide by x and you'll hvae a linear equation

OpenStudy (anonymous):

y' - y/x = x^2 :/ nothing changes

OpenStudy (anonymous):

solve the homogenous equation \[y'-\frac{1}{x}y=0\]

OpenStudy (anonymous):

are you sure about that it is a part of variation of parameters method

OpenStudy (anonymous):

i do not see like that

OpenStudy (anonymous):

well the first part in my book uses an auxilary equation however since you do not have the x variable you need to find it out another way.... variation of parameter has to deal with the particular solution if i'm not mistaken

OpenStudy (anonymous):

hmm i guess you are right i do not have much information about that

OpenStudy (anonymous):

yeah you have to have a way to find the complementary function

OpenStudy (anonymous):

y = (x^3/2) + cx

OpenStudy (anonymous):

one way that you might have learn is to use linear to find one solution and then using that solution to find a y_2

OpenStudy (anonymous):

@Outkast3r09 i found the solution of h.eq but is it right ?

OpenStudy (anonymous):

hold on i find this weird because usually variation of parameters is ued int higher order equations

OpenStudy (anonymous):

yeah you are right

OpenStudy (anonymous):

are you sure that's what your teacher wrote?

OpenStudy (anonymous):

in the book :)

OpenStudy (anonymous):

if my eyes are not damaged , it is true :D

OpenStudy (dumbcow):

since this is a first order DE, just find the integrating factor

OpenStudy (anonymous):

it is 1/x

OpenStudy (dumbcow):

yes, multiply 1/x to every term

OpenStudy (anonymous):

that's what i'm saying dumbcow? you'll find the general solution

OpenStudy (anonymous):

yeah i do that and it was y = (x^3/2) + cx

OpenStudy (anonymous):

if you do the non homogenous however reduction formula is used to find another general solution but you cannot use it... if you solve the nonhomog linear you'l get y=yc+yp which you do not need to use variation of parameters on it

OpenStudy (anonymous):

yeah it is easy but book says the variation of parameter :(

OpenStudy (anonymous):

is it a darker prime... i'm looking at another site and it has a darker prime mark to show y''

OpenStudy (anonymous):

personally i believe this is a typo i'd ask my teacher about it

OpenStudy (anonymous):

o-oo i do not think so i even do not know what darker prime is

OpenStudy (anonymous):

hmm i have to give it tomorrow :/